# Benchmarking Zero-Shot Facial Emotion Annotation with Large Language Models: A Multi-Class and Multi-Frame Approach in DailyLife

### v1

@@ 4.4. Performance of Different Strategies: Insights from Confusion Matrices
The presented confusion matrices (in Fig. 3) compare the performance of various classification strategies (A1, A2, B1, B2, C1, C2, D1, D2) across different tasks involving emotion and sentiment recognition. Each matrix visualizes the true labels versus the predicted labels, providing insights into the model’s accuracy, strengths, and areas requiring improvement. Strategies A1, B1, C1, and D1 are evaluated on their ability to classify seven distinct emotional states, including “Angry,” “Happy,” “Neutral,” and “Sad.” The diagonal entries reflect correct classifications, while off-diagonal values indicate confusion between emotions. For example, significant misclassifications are observed between “Neutral” and “Happy” in some strategies, highlighting challenges in distinguishing subtle emotional variations. Strategies A2, B2, C2, and D2 focus on sentiment classification into three categories: “Negative,” “Neutral,” and “Positive.” While these strategies generally achieve high accuracy for the “Negative” category, frequent confusion between “Neutral” and “Positive” suggests a need for improved sensitivity to nuanced sentiment expressions.
Different strategies exhibit similar distribution patterns, indicating that the zero-shot LLM annotation approach introduces a certain level of ambiguity in labeling the aforementioned emotion categories. This ambiguity often leads to confusion between certain categories, while labels for negative emotions are generally more accurate. This observation suggests potential opportunities for emotion recognition tasks in specific contexts, where leveraging the strengths of zero-shot LLM annotation could enhance performance despite its inherent limitations.
@@ 4.5.1. Baseline Comparison (Random Guessing)
To contextualize the performance of our annotation strategies, we compared them against baseline accuracy levels derived from random guessing. In the seven-class taxonomy, random guessing would yield an expected accuracy of approximately 14.3%, while in the three-class taxonomy, the expected accuracy is around 33.3%. Our results demonstrate that all proposed strategies significantly surpass these baseline levels. Specifically, in the seven-class taxonomy, the best-performing strategies (C1 and D1) achieved an accuracy of 46%, more than three times the baseline. Similarly, in the three-class taxonomy, Strategies B2, C2, and D2 reached an accuracy of 65%, nearly doubling the random guessing baseline. This substantial improvement underscores the effectiveness of our aggregation and integration methods in enhancing annotation accuracy within zero-shot classification tasks using large language models.
4.5.2. Baseline Comparison (Trained Models)
@@
Our best-performing strategy within the seven-class taxonomy, Strategy D1 (Multi-Frame Integration), achieved a WAR of 46%, closely approaching the performance of the VGG13-LSTM (46.07% WAR) and Two VGG13-LSTM (46.92% WAR) models—both of which represent the top-performing baseline methods. Moreover, Strategy D1 significantly surpasses the average WAR of the baseline models, which stands at approximately 38.98%.
In terms of UAR, Strategy D1 outperforms all baseline models, achieving the highest recall across all classes without being influenced by class imbalance. This indicates that our strategy not only excels in overall weighted performance but also ensures equitable recognition of all emotion categories, including those that are less frequent in dataset.
@@
Average 38.98 26.75
This Study Zero-Shot LLMα 46.00 35.00
@@
Table 3. Comparison of Weighted Average Recall (WAR) and Unweighted Average Recall (UAR) between baseline architectures (trained from scratch) and the LLM-based zero-shot method (GPT-4o-mini) on the FERV39k DailyLife subset (Wang et al., 2022).
@@ 5. Cost-Efficiency and Scalability
Besides performing close to or outperforming baseline models in performance, compared to traditional supervised models, GPT-4o-mini-based annotation strategies have significant advantages in cost-effectiveness and scalability. Strategy D1 reduces operational costs by minimizing the number of API requests and decreasing token lengths through preprocessing. This cost-effective approach ensures that large-scale annotation tasks remain financially feasible. Furthermore, our zero-shot annotation approach leverages the capabilities of LLMs without necessitating task-specific training, allowing for rapid deployment and adaptation to various annotation tasks with minimal additional resources.
While some baseline models, such as Two VGG13-LSTM, exhibit marginally higher WAR scores, our annotation strategy D1 achieves comparable performance levels with enhanced cost and operational efficiencies. This underscores the effectiveness of aggregation techniques and temporal context integration in zero-shot annotation tasks using LLMs, presenting a viable alternative to traditional supervised models, especially in scenarios constrained by budget and computational resources.
6. Model Cost Considerations
In the initial phase of our study, we tested full-frame rate annotation, labeling all 25 images for each second. However, given the task volume, this approach was financially unsustainable, with API costs reaching approximately \$100 for annotating $\sim$11,000 images among this dataset. To mitigate costs, we adopted a strategy of selecting five key frames (the first, Q1, middle, Q3, and final frames) from each video segment. Additionally, we merged these five frames into a single input, significantly reducing token usage. Although we did not perform frame-by-frame annotation at full frame rates, our results remain valuable for tasks that are not highly sensitive to high frame rates.
7. Ethical Considerations
LLMs, as a “technological revolution,” have brought unprecedented opportunities, driving paradigm shifts in tasks including, but not limited to, those outlined in this study. While our research explores the technical potential of LLMs in emotion annotation tasks, it is essential not to overlook the associated ethical considerations. We hope to take this opportunity to remind researchers employing this method to use it responsibly and thoughtfully, particularly under the concept of superalignment.
@@ 8. Conclusion with Future Work
This study demonstrates the feasibility of using LLMs for automated emotion annotation in facial images through zero-shot classification. By exploring various annotation strategies, we identified the potential of LLMs to achieve competitive performance, particularly in tasks involving ternary classification of emotions. Strategies that integrate multiple frames or aggregate annotations through majority voting significantly enhance the reliability of emotion recognition, offering a promising alternative to traditional supervised methods.
While LLMs exhibit inherent ambiguity in distinguishing closely related emotion categories, particularly within the seven-class taxonomy, they achieve higher accuracy in simpler classification tasks. This highlights their utility in scenarios where efficiency and scalability are prioritized over fine-grained classification precision. Moreover, the cost-effective nature of zero-shot LLM annotation enables large-scale deployment, reducing the reliance on human annotators and minimizing operational costs.
Our findings underscore the importance of leveraging aggregation techniques, temporal context, and task simplification to maximize the potential of LLMs in emotion annotation. Future work should focus on fine-tuning multimodal LLMs for emotion recognition, addressing ambiguities in classification, and expanding their application in real-world multimodal environments, such as driver attention detection, live streaming platform moderation, and health management systems. This research provides a foundation for advancing automated annotation techniques, fostering innovation in human-computer interaction and affective computing domains.

### v2

@@ 4.4. Performance of Different Strategies: Insights from Confusion Matrices
The presented confusion matrices compare the performance of various classification strategies (A1, A2, B1, B2, C1, C2, D1, D2) across different tasks involving emotion and sentiment recognition. Each matrix visualizes the true labels versus the predicted labels, providing insights into the model’s accuracy, strengths, and areas requiring improvement. Strategies A1, B1, C1, and D1 are evaluated on their ability to classify seven distinct emotional states, including “Angry,” “Happy,” “Neutral,” and “Sad.” The diagonal entries reflect correct classifications, while off-diagonal values indicate confusion between emotions. For example, significant misclassifications are observed between “Neutral” and “Happy” in some strategies, highlighting challenges in distinguishing subtle emotional variations. Strategies A2, B2, C2, and D2 focus on sentiment classification into three categories: “Negative,” “Neutral,” and “Positive.” While these strategies generally achieve high accuracy for the “Negative” category, frequent confusion between “Neutral” and “Positive” suggests a need for improved sensitivity to nuanced sentiment expressions.
Different strategies exhibit similar distribution patterns, indicating that the zero-shot LMM annotation approach introduces a certain level of ambiguity in labeling the aforementioned emotion categories. This ambiguity often leads to confusion between certain categories, while labels for negative emotions are generally more accurate. This observation suggests potential opportunities for emotion recognition tasks in specific contexts, where leveraging the strengths of zero-shot LMM annotation could enhance performance despite its inherent limitations.
@@ 4.5.1. Baseline Comparison (Random Guessing)
To contextualize the performance of our annotation strategies, we compared them against baseline theoretical chance levels derived from random guessing. In the seven-class taxonomy, random guessing would yield an expected accuracy of approximately 14.3%, while in the three-class taxonomy, the expected accuracy is around 33.3%. Our results demonstrate that all proposed strategies significantly surpass these baseline levels. Specifically, in the seven-class taxonomy, the best-performing strategies (C1 and D1) achieved an accuracy of 46%, more than three times the baseline. Similarly, in the three-class taxonomy, Strategies B2, C2, and D2 reached an accuracy of 65%, nearly doubling the random guessing baseline. This substantial improvement underscores the effectiveness of our aggregation and integration methods in enhancing annotation accuracy within zero-shot classification tasks using large language models.
4.5.2. Baseline Comparison (Pre-trained Models and Zero-shot Methods)
@@
AEN (Lee et al., 2023) 47.88 38.18
Average* 39.97 27.87
CLIP (Radford et al., 2021)β 17.10 20.99
FaRL (Zheng et al., 2022)β 25.65 21.67
EmoCLIP (Foteinopoulou and Patras, 2024)β 35.30 26.73
Averageβ 26.02 23.13
FineCLIPER (Chen et al., 2024)γ 48.63 40.79
Average 38.39 27.81
This Study Zero-Shot LMMα 46.00 35.00
*Pre-trained
@@
βThe result of zero-shot method
γThe result of zero-shot method with label augmentation
Table 7. Comparison of Weighted Average Recall (WAR) and Unweighted Average Recall (UAR) between baseline architectures (trained from scratch) and the LMM-based zero-shot method (GPT-4o-mini) on the FERV39k DailyLife subset (Wang et al., 2022).
Our best-performing strategy within the seven-class taxonomy, Strategy D1 (Multi-Frame Integration), achieved a WAR of 46%, which closely rivals top-performing pre-trained baseline methods such as AEN (47.88% WAR), VGG13-LSTM (46.07% WAR) and Two VGG13-LSTM (46.92% WAR), while significantly exceeding the baseline average of approximately 38.98% WAR. Moreover, Strategy D1 significantly surpasses the average WAR of the baseline models, which stands at approximately 38.98%. Compared to other zero-shot methods, our Strategy D1’s result are higher than the results of zero-shot methods. Although FineCLIPER (48.63% WAR and 40.79% UAR) achieves better performance, it utilizes additional label augmentation, which we have not employed. We suggest that future research could explore a combination of FineCLIPER’s and our methods, aiming to further save computational resources while still improving recognition results.
@@ 5. Cost-Efficiency and Scalability
Besides performing close to or outperforming baseline models in performance, compared to traditional supervised models, GPT-4o-mini-based annotation strategies have significant advantages in cost-effectiveness and scalability. Strategy D1 reduces operational costs by minimizing the number of API requests and decreasing token lengths through preprocessing. This cost-effective approach ensures that large-scale annotation tasks remain financially feasible. Furthermore, our zero-shot annotation approach leverages the capabilities of LMMs without necessitating task-specific training, allowing for rapid deployment and adaptation to various annotation tasks with minimal additional resources.
While some baseline models, such as Two VGG13-LSTM, exhibit marginally higher WAR scores, our annotation strategy D1 achieves comparable performance levels with enhanced cost and operational efficiencies. This underscores the effectiveness of aggregation techniques and temporal context integration in zero-shot annotation tasks using LMMs, presenting a viable alternative to traditional supervised models, especially in scenarios constrained by budget and computational resources.
@@ 6. Conclusion with Future Work
This study demonstrates the feasibility of using LMMs for automated emotion annotation in facial images through zero-shot classification. By exploring various annotation strategies, we identified the potential of LMMs to achieve competitive performance, particularly in tasks involving ternary classification of emotions. Strategies that integrate multiple frames or aggregate annotations through majority voting significantly enhance the reliability of emotion recognition, offering a promising alternative to traditional supervised methods.
While LMMs exhibit inherent ambiguity in distinguishing closely related emotion categories, particularly within the seven-class taxonomy, they achieve higher accuracy in simpler classification tasks. This highlights their utility in scenarios where efficiency and scalability are prioritized over fine-grained classification precision. Moreover, the cost-effective nature of zero-shot LMM annotation enables large-scale deployment, reducing the reliance on human annotators and minimizing operational costs.
Our findings underscore the importance of leveraging aggregation techniques, temporal context, and task simplification to maximize the potential of LMMs in emotion annotation. Future work should focus on fine-tuning multimodal LMMs for emotion recognition, addressing ambiguities in classification, and expanding their application in real-world multimodal environments, such as driver attention detection, live streaming platform moderation, and health management systems. This research provides a foundation for advancing automated annotation techniques, fostering innovation in human-computer interaction and affective computing domains.
Safe and Responsible Innovation Statement
We hope to take this opportunity to remind researchers employing this method to use it responsibly and thoughtfully, particularly under the concept of superalignment. It is also important to note that LLMs may hallucinate in the label space, potentially introducing inaccurate or spurious outputs that could mislead downstream processes. In this paper, our tests were carried out by calling the API, and under the privacy policy, none of these data will be shared or used for training purposes. In our study, the testing was carried out via API calls; under the privacy policy, this data will neither be shared nor used for training purposes. We also recommend experimenting with on-premise models in the future to eliminate the risk of data leaks. Beyond that, when prototyping responsibly around this capability, it may be wise to target specific tasks and scenarios first rather than rolling it out directly to a wider market.
Acknowledgements.
We thank the anonymous reviewers for their valuable feedback. We also acknowledge the authors of the FERV39k dataset for their contributions and for granting permission for its non-commercial use. This work was supported by the Beijing Natural Science Foundation-Youth Project (Grant No.4254082).